At one high school, students can run the 100-yard dash in an average of 15.2 seconds with a standard deviation of .9 seconds. The times are very closely approximated by a normal curve. Find the percent of times that are: greater than 16.1 seconds.
Answer by stanbon(48568)

At one high school, students can run the 100-yard dash in an average of 15.2 seconds with a standard deviation of .9 seconds. The times are very closely approximated by a normal curve. Find the percent of times that are: greater than 16.1 seconds.
--------------------
z(16.1) = (16.1-15.2)/0.9 = 1
----------
P(x>16.1)= P(z>1) = 0.1587.. = 15.87%
=================
Cheers,
Stan H.